OpenStudy (rock_mit182):

tough one : Ammonium perchlorate (NH4ClO4) is a solid rocket fuel used in space shuttles. When heated above 200°C, it decomposes to a variety of gaseous products, of which the most important are N2, Cl2, O2, and water vapor. (c) Calculate the mole fraction of chlorine and its partial pressure in the mixture of gases produced.

OpenStudy (rock_mit182):

(b) The sudden appearance of hot gaseous products in a small initial volume leads to rapid increases in pressure and temperature, which give the rocket its thrust. What total pressure of gas would be produced at 800°C by igniting 7.00 × 105 kg NH4ClO4 (a typical charge of the booster rockets in the space shuttle) and allowing it to expand to fill a volume of 6400 m3 (6.40 × 106 L)? Use the ideal gas law.

OpenStudy (anonymous):

8n?

OpenStudy (rock_mit182):

\[PV = nR DeltaT\]

OpenStudy (rock_mit182):

C is no hard, but B ?

OpenStudy (rock_mit182):

altought id like someone explain me conceptually both

OpenStudy (ipwnbunnies):

B is simple, just a plug n chug into the ideal gas formula. But piggybacking off apurba, can you write the balance equation clearer?

OpenStudy (ipwnbunnies):

Where's 8n coming from? What is that?

OpenStudy (rock_mit182):

oh man my bad,wait I'll rewrite it

OpenStudy (rock_mit182):

2 NH4ClO4(s) ---> N2(g) + Cl2(g) + 2 O2(g) + 4 H2O(g)

OpenStudy (ipwnbunnies):

Ok, so what we're trying to do is find the pressure of the gas, which we can find with PV = nRT. Rearranging it: P = nRT/V

OpenStudy (ipwnbunnies):

We do have the #s, but they're not in correct form where we can plug it in, assuming we're using gas constant R = 0.08206 (L*atm)/(mol*K)

OpenStudy (rock_mit182):

so we have to convert

OpenStudy (ipwnbunnies):

First, n, which is the # of moles of gas: We have the mass of gas: 105 kg = 105e3 g. Find the molar mass of NH4ClO4 in g/mol and divide the 2 to get the # of moles of gas.

OpenStudy (ipwnbunnies):

Next, T: We need to get the temperature in Kelvin instead of Celsius. Temp(K) = Temp(C) + 273.15

OpenStudy (rock_mit182):

X mol NH4ClO4 = 1 mol NH4ClO4* 105 * 10^3 g/ # of grams in NH4ClO4

OpenStudy (ipwnbunnies):

Finally, volume: We need volume is Liters. You did write 6400 m^3 = 6.4e6 L, which is correct.

OpenStudy (rock_mit182):

well i think were harder. man and the (c) can you explain me what does it mean conceptually

OpenStudy (rock_mit182):

i mean the first one

OpenStudy (ipwnbunnies):

For c, mole fraction for Chlorine is the # of moles of Cl / # of total moles.

OpenStudy (ipwnbunnies):

We found the # of total moles in Part B. Use stoichiometry from the balanced eq. to find the # of moles of Chlorine. mol Cl/total mol would be the mole fraction, which is X symbolically I believe.

OpenStudy (ipwnbunnies):

To find the partial pressure of Cl, you would need to multiply the mole fraction of Cl and the total pressure of the gas, which is the answer of Part B.

OpenStudy (rock_mit182):

thanks a lot brb